OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x.

OpenStudy (tkhunny):

"Confirm" -- This is nice. Just follow the instructions. That's what "confirm" is all about.

OpenStudy (anonymous):

\[f(x) = \frac{ x-7 }{ x+3 } ; g(x)=\frac{ -3x-7 }{ x-1 }\]

OpenStudy (tkhunny):

Well? If it were f(3), what would you do?

OpenStudy (anonymous):

plug in 3 to every (x)

OpenStudy (tkhunny):

Do that with g(x) and you will have f(g(x)). It's just what the notation means.

OpenStudy (anonymous):

so for every (x) in f(x) i plug in -3x - 7 / x - 1 ?

OpenStudy (tkhunny):

You have it. Let's see it.

OpenStudy (anonymous):

\[\frac{ -3x - 7/ x - 1 - 7}{ -3x - 7/ x - 1 +3 } \]

OpenStudy (tkhunny):

Horrid notation. Please use parentheses to clarify intent. f(x) = (x-7)/(x+3) g(x) = (-3x-7)/(x-1) f(g(x)) = ([(-3x-7)/(x-1)]-7)/([(-3x-7)/(x-1)]+3) It's not pretty, but it's complete and accurate. Now, for your best algebra skills.

OpenStudy (anonymous):

my apologies. can i cross out ( -3x - 7 / x - 1 ) ?

OpenStudy (anonymous):

or do i multiply ( x - 1 ) for numerator and denominator?

OpenStudy (tkhunny):

You simplify. Let's take the numerator. \(\dfrac{-3x-7}{x-1} - 7 = \dfrac{-3x-7}{x-1} - \dfrac{7(x-1)}{x-1} = \dfrac{(-3x-7)-7(x-1)}{x-1}\) Keep going. We're just adding fractions.

OpenStudy (anonymous):

for the numerator do i distribute?

OpenStudy (tkhunny):

You do what it takes to simplify it. If the Distributive Property is appropriate, then do that. \(= \dfrac{-3x - 7 - 7x + 1}{x-1}\) One step at a time.

OpenStudy (anonymous):

from that i got \[\frac{ -10x - 6 }{ x - 1 }\]

OpenStudy (tkhunny):

Okay, now tackle the denominator.

OpenStudy (anonymous):

could you start me off so i know where to begin?

OpenStudy (anonymous):

be prepared to do a raft of algebra ready?

OpenStudy (anonymous):

very ready!

OpenStudy (anonymous):

first we compute \[f(g(x)) = x\]

OpenStudy (anonymous):

\[f(g(x))=f(\frac{ -3x-7 }{ x-1 })\]

OpenStudy (anonymous):

now we are going to replace all \(x\) in \(f(x)\) by \(\frac{ -3x-7 }{ x-1 }\) that is actually very easy for me to do here, by cutting and pasting

OpenStudy (tkhunny):

You already have the denominator. You MUST show some algebra. ([(-3x-7)/(x-1)]+3) = \(\dfrac{-3x-7}{x-1}+3\) Go!

OpenStudy (anonymous):

\[f(x) = \frac{ x-7 }{ x+3 }\]so \[f(g(x))= \frac{ \frac{ -3x-7 }{ x-1 }-7 }{ \frac{ -3x-7 }{ x-1 }+3 }\]

OpenStudy (tkhunny):

THERE'S that denominator!

OpenStudy (anonymous):

to get rid of that annoying compound fraction, multiply the numerator and denominator by \(x-1\) (carefully using parentheses)

OpenStudy (anonymous):

then multiply out, combine like terms since you know the answer will just be \(x\) you should expect an orgy of cancellation at the last couple steps

OpenStudy (anonymous):

you need the first step?

OpenStudy (solomonzelman):

(I would rather just find the inverse, without doing f(g(x))=x and vv, but ... \(-:(\) )

OpenStudy (anonymous):

uh, i got \[\frac{ -10x - 6 }{ 8 }\]

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

\[ \frac{ \frac{ -3x-7 }{ x-1 }-7 }{ \frac{ -3x-7 }{ x-1 }+3 }\] the \(x-1\) will cancel top and bottom to get \[\frac{-3x-7-7(x-1)}{-3x-7+3(x-1)}\]

OpenStudy (anonymous):

now multiply out using the almighty distributive law then combine like terms

OpenStudy (anonymous):

then \[\frac{ -3x - 7 - 7x + 1 }{ -3x - 7 + 3x - 1 }\]

OpenStudy (anonymous):

forgot that distributive law already huh?

OpenStudy (anonymous):

distribute the \(-7\) up top and the \(3\) below

OpenStudy (anonymous):

starting here \[\frac{-3x-7-7(x-1)}{-3x-7+3(x-1)}\]

OpenStudy (anonymous):

(-3x - 7 ) ( -7x + 1) (-3x - 7 ) ( 3x - 1)

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

what can you get with \[-7(x-1)\]when you distribute?

OpenStudy (anonymous):

-7x + 7 OMG -.- im so disappointed with myself right now.

OpenStudy (anonymous):

ok so lets back up to \[\frac{-3x-7-7(x-1)}{-3x-7+3(x-1)}\] and see what we get when we remove the parentheses by "we" i mean "you"

OpenStudy (anonymous):

i got, -10x / - 10 which would equal to (x) so we have f(g(x)) out of the way, i started g(f(x)), coud you guide me with that to please?

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

ok sure lets start just as before with an annoying compound fraction but try it yourself first, it is going to work almost exactly like this one

OpenStudy (anonymous):

\[(-3 (\frac{ x-7 }{ x+3 }) - 7 (numerator) ; (\frac{ x-7 }{ x+3}) - 1 (denominator)\]

OpenStudy (anonymous):

let me try to write it

OpenStudy (anonymous):

yeah actually that looks good now multiply top and bottom but this time by \(x+3\) instead of \(x-1\)

OpenStudy (anonymous):

\[ g(x)=\frac{ -3x-7 }{ x-1 }\] \[ g(f(x))=\frac{ -3 \frac{ x-7 }{ x+3 } -7 }{ \frac{ x-7 }{ x+3 } -1 }\]

OpenStudy (anonymous):

should have used parentheses around the first term in the top

OpenStudy (anonymous):

multiply by \(x+3\) top and bottom what do you get before distributing etc

OpenStudy (anonymous):

this might be trickier because of the \(-3\) up top so be careful with parentheses

OpenStudy (anonymous):

-3(x-7) - 7(x+3) / (x-7) - 1(x+3) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now get rid of the parentheses

OpenStudy (anonymous):

you should of course just be left with some x up top and a number in the bottom that will cancel with it

OpenStudy (anonymous):

-10x / - 10 like before we get (x) for g(f(x)). whew! long problem lol. but thank you sooooo much for all of your time & help!! i wish i could give you more than just a medal & a fan!

OpenStudy (anonymous):

glad to help, hope you learned something (at least learned how to do these!) who studies math in late august?